# Navigating Through Paper Flood: Advancing LLM-based Paper Evaluation through Domain-Aware Retrieval and Latent Reasoning

### v1

@@ Introduction
Technically, the paper evaluation task aims to analyze the paper features to assess research quality from various dimensions, such as academic impact (Xia, Li, and Li 2022; Zhao et al. 2024) and overall quality (Lin et al. 2023). Existing studies mainly rely on traditional neural models or Large Language Models (LLMs) for this task:
  • Traditional methods utilize neural models, such as Multi-Layer Perceptrons (MLPs), or Long-Short Term Memory networks (LSTM), to evaluate research papers based on predefined features, including structural indicators like paper length and reference count (Vergoulis et al. 2020; Ruan et al. 2020), as well as textual patterns (Ma et al. 2021; Yang et al. 2018). However, these methods often overlook the semantic content of papers (e.g., abstract and main text), which is essential for accurate evaluation, ultimately leading to unsatisfactory performance.
  • LLM-based methods leverage rich textual information (e.g., title, abstract, and main text) to learn informative paper representations and employ a scoring module to produce evaluation scores. Empowered by the advanced semantic understanding capabilities of LLMs, these methods demonstrate strong potential in capturing the technical soundness of research papers, yielding more accurate evaluation results (Lu et al. 2024; Liu et al. 2025a; Zhao et al. 2025; de Winter 2024).
@@
  • Reasoning-enhanced paper evaluation. Evaluating research papers requires a deep understanding of complex motivations and nuanced methodologies. This challenge is further intensified by the need to compare concurrently retrieved work. This motivates us to stimulate the reasoning mechanism of LLMs to support deep comprehension, precise comparison, and fair evaluation. While chain-of-thought reasoning (Wei et al. 2022) offers interpretable intermediate steps, it typically requires annotated reasoning paths for supervision (Weng et al. 2023), which are scarce in paper evaluation scenarios. In contrast, latent reasoning (Hao et al. 2024) enables implicit multi-step reasoning within the hidden representations of LLMs, eliminating the need for explicit annotations. More importantly, LLM-based paper evaluation focuses on learning more informative paper representations to enhance evaluation accuracy, which aligns naturally with the latent reasoning paradigm, seamlessly integrating reasoning directly at the representation level. As such, we incorporate latent reasoning into PaperEval for comprehensive representation learning.
Despite the significant potential of latent reasoning, an effective optimization strategy is essential to guide the reasoning process toward the ultimate ranking goal of paper evaluation. Specifically, the paper evaluation task focuses on comparing and identifying valuable work within a large collection, with a primary emphasis on relative ranking rather than absolute scoring. However, learning accurate rankings is inherently more challenging, as even minor prediction errors may cause substantial shifts in the ranking positions. To address this, we propose a progressive ranking optimization strategy, which encourages the latent reasoning to progressively improve relative ranking. In particular, at each reasoning step, we compute the temperature-controlled softmax over the predicted scores of a batch of papers, which is then aligned with the ground-truth order using a listwise ranking loss. To gradually refine the LLM’s ranking, we progressively decrease the temperature during latent reasoning, making the predicted distributions increasingly sharper and more sensitive to ranking errors. This progressive refinement encourages the LLM to iteratively produce more confident and discriminative rankings within each training batch. By learning to distinguish fine-grained differences among batch samples, the LLM enhances ranking reasoning capabilities, which can naturally generalize to global ranking across the entire dataset, as theoretically supported by (Lan et al. 2009).
@@
Paper evaluation aims to assess a paper’s quality or predict its academic impact. From the quality perspective, a central task is paper rating — predicting whether a paper will be accepted by peer review committees (Lin et al. 2023). Existing methods fall into three main categories. The first leverages neural architectures like CNNs and attention-based models to capture local and global textual interactions (Yang et al. 2018; Deng et al. 2020). The second extracts metadata features and employs traditional models such as random forests (Wang et al. 2024). The third directly encodes the textual content using pretrained models like BERT (Devlin et al. 2019), as in recent work (Xue et al. 2023; Liu et al. 2025a). For evaluating academic impact, prior studies focus on predicting citation counts, citation levels, or other derived impact metrics (Zhao et al. 2024), which can similarly be grouped into three strategies. Metadata-based approaches use handcrafted or extracted features with classical models like MLPs or decision trees (Wang, Yu, and Yu 2011; Qiu and Han 2024; Ruan et al. 2020; Zhang and Wu 2024). Graph-based methods model early citation dynamics using citation graphs and apply graph neural networks for future trend prediction (Yan et al. 2024; He et al. 2023; Li et al. 2023; Jiang, Koch, and Sun 2021). LLM-based approaches either prompt models with a paper’s title and abstract to generate citation scores (de Winter 2024), or map dense representations to impact-related metrics (Zhao et al. 2025). Despite these advances, accurate and fine-grained paper evaluation remains challenging. In this work, we explore how integrating retrieval techniques with the reasoning capabilities of LLMs can address this challenge.
@@
Unlike Chain-of-Thought (CoT) reasoning (Wei et al. 2022; Su et al. 2025; Carrow et al. 2025), which relies on explicitly generated intermediate steps, latent reasoning performs inference directly within a model’s hidden representations (Hao et al. 2024; Biran et al. 2024). This implicit approach has gained momentum in LLM-based recommendation (Tang et al. 2025; Shen et al. 2025a, b; Liu et al. 2025b) and retrieval tasks (Ji et al. 2025), as it avoids the need for annotated reasoning traces while enabling richer, more informative representations. However, a key challenge remains: how to effectively supervise the latent reasoning process. On the one hand, it is essential to ensure that the reasoning process unfolds progressively toward the correct output, allowing the model to refine its prediction step by step (Tang et al. 2025; Liu et al. 2025b; Ji et al. 2025). On the other hand, models must avoid degenerate reasoning, where the hidden states prematurely converge and hinder iterative refinement. For instance, Tang et al. (Tang et al. 2025) introduce a loss that encourages representational diversity across reasoning steps to mitigate this issue. Despite these efforts, ensuring that the model consistently refines its predictions toward accurate outcomes remains challenging. In this work, we propose a progressive optimization strategy that progressively adjusts the softmax temperature and incorporates a ranking-based loss. This design explicitly guides the latent reasoning process toward better alignment with ground-truth evaluation targets.
@@
To equip LLMs with up-to-date domain knowledge for more accurate assessment of the novelty and contributions of each target paper, we introduce a domain-aware retrieval module (depicted at the top of Figure 1). Specifically, given the title and abstract of research papers, we first employ ChatGPT (Achiam et al. 2023) to generate representative topic keyphrases, which are then encoded into topic embeddings using the CLIP text encoder (Radford et al. 2021). To identify relevant work for each target paper $p_{i}$, we first compute the cosine similarity between the target topic embedding and those of all other papers in the corpus, where papers with similarity exceeding a predefined threshold $\gamma$ are retained as candidates. Considering the rapidly evolving nature of many research fields, we further filter these candidates by selecting only concurrent relevant papers, whose publication dates are closest to the target paper $p_{i}$. The resulting set, denoted as $\mathcal{R}_{i}$, contains at most $k$ papers and serves as the domain-aware reference set, providing essential contextual background to help the LLM more accurately assess the target paper’s position within the current research landscape.
@@
Inspired by ListMLE (Xia et al. 2008), which learns to predict rankings by maximizing the likelihood of the ground-truth order, we adapt it to the paper evaluation scenario, encouraging the predicted scores to yield a ranking consistent with the ground-truth.
@@
(1) NAID (Zhao et al. 2025): This dataset is derived from arXiv. Each paper includes its title, abstract, some metadata (e.g., paper length, number of references), and an impact score that quantifies its relative citation rank within the same domain and publication period, serving as an indicator of the paper’s future impact in its field. The dataset contains 11,118 papers in the training set and 1,237 papers in the test set.
@@
  • Traditional methods: 1) MLP-based (Ruan et al. 2020) method uses metadata of the target paper as input to an MLP to predict the evaluation score. We do not evaluate this method on the ICLR dataset due to the lack of metadata. 2) LSTM-based (Ma et al. 2021) method encodes the abstract of each paper into a sequence representation and applies an LSTM network to predict the target score.
  • LLM-based methods: 3) GPT-part (de Winter 2024) method prompts ChatGPT to predict a paper’s score based solely on its title and abstract. We use GPT-4o (Hurst et al. 2024) as the underlying model. 4) GPT-all (Lu et al. 2024) approach treats the LLM as a reviewer, prompting it to read the entire paper and generate a full review, including an overall quality score. For cost considerations, we choose GPT-4o-mini as our base model. 5) SciBERT (Beltagy, Lo, and Cohan 2019) method is a BERT-based model pretrained on scientific text. We fine-tune it with a simple regression module to perform paper evaluation. 6) NAIP (Zhao et al. 2025) method uses LLaMA3-Smaug (Pal et al. 2024) as the backbone LLM. An additional regression module is applied to the output embedding to generate the final score.
@@
  • Top-K ranking quality: Following (Zhao et al. 2025), we use NDCG@{10,20} to measure the ranking quality of the top-K recommended papers.
  • Overall ranking consistency: Following (Ng and Abrecht 2015), we employ Spearman’s rho and Kendall’s tau to assess how well the predicted rankings align with the ground-truth rankings.
@@
  • Pair-wise ranking: Inspired by RankNet (Burges et al. 2005), we design a temperature-controlled pair-wise ranking loss to examine whether pair-wise supervision is more suitable for PaperEval than list-wise ranking.
@@
  • Score regression: As a control setting, similar to (Zhao et al. 2025), we remove the ranking-based objective and directly apply an MSE loss between the final predicted score and the ground-truth label. This setup allows us to examine whether ranking is a more effective learning signal than direct score supervision.
@@ Appendix A Supplementary Material
Code and Datasetshttps://anonymous.4open.science/r/PaperEval-85D7
@@
Following NAIP (Zhao et al. 2025), we adopt LLaMA3-Smaug (Pal et al. 2024) as the backbone of PaperEval for fair comparison. We set the number of reasoning steps $m$ to 8 and the number of retrieved reference papers $k$ to 2. We set the temperature $\tau_{\text{max}}$ to $1.0$ and $\tau_{\text{min}}$ to $0.1$. We train the model for 5 epochs on both the NAID and ICLR datasets with a shared learning rate of $5.0\times 10^{-5}$, and select the checkpoint with the best validation performance for evaluation. All experiments are conducted on a single NVIDIA A40 GPU.
@@
Inspired by (Yang et al. 2020), we also design a distribution-based similarity/ListNet (Cao et al. 2007) loss to encourage the predicted score distribution to match the ground-truth score distribution. Specifically, we convert both the predicted scores and the ground-truth scores into soft probability distributions using a temperature-controlled softmax. The temperature varies with the reasoning step to enable coarse-to-fine comparison. The ground-truth distribution at step $j$ is computed as:
@@
We also try the RankCosine ranking loss (Qin et al. 2008) in PaperEval. RankCosine treats the predicted and ground-truth score lists as vectors and maximizes their cosine similarity, encouraging the predicted ranking to align closely with the target. In addition, we introduce temperature control to progressively sharpen the predicted vector during reasoning steps as follows:
@@
The Approximate NDCG (Qin, Liu, and Li 2010) loss is a differentiable surrogate for the standard NDCG metric, enabling its direct optimization in learning-to-rank models. It overcomes the non-differentiability of the ranking operation by replacing the discrete integer rank of an item with a continuous “soft rank” $\hat{\pi}_{i}$. This soft rank is calculated from pairwise comparisons of item scores using a sigmoid function $\sigma$ scaled by a temperature parameter $\tau$.
@@
We adopt a pair-wise ranking loss inspired by RankNet (Burges et al. 2005) to guide the model in learning relative paper quality comparisons. Specifically, at each evaluation step $t\in\{1,2,\cdots,m\}$ we focus on all pairs of papers $(i,j)$ where paper $i$ is preferred over paper $j$, means $s_{i}>s_{j}$. The model predicts scalar scores $s_{i}^{(t)}$ and $s_{j}^{(t)}$ for the two papers at step $t$, and the loss encourages $s_{i}^{(t)}$ to be higher than $s_{j}^{(t)}$ . The pair-wise loss at step $t$ is defined as:
@@
We conduct all loss design experiments on the NAID dataset. From the results in Table 4, we make two key observations: 1) ListMLE achieves the best performance among all loss functions, as its permutation probability modeling directly reflects the ground-truth ranking, making it particularly well-suited to our training objective. This also supports the theoretical analysis in (Lan et al. 2009). 2) Most other listwise and pairwise losses fail to outperform simple MSE, indicating that effective ranking supervision remains challenging compared to direct score prediction.
@@
Llama* (Pal et al. 2024) 8B 0.9589 0.9521 0.4953 0.3438 0.7784 0.7386 0.3276 0.2285
Llama (Grattafiori et al. 2024) 8B 0.9724 0.9038 0.4895 0.3405 0.7276 0.7306 0.3004 0.2091
Qwen (Qwen et al. 2025) 7B 0.9691 0.9314 0.4950 0.3446 0.7698 0.7554 0.3332 0.2320
Mistral (Jiang et al. 2023) 7B 0.8260 0.8355 0.4267 0.2944 0.7055 0.7041 0.3055 0.2141
Phi (Abdin et al. 2024) 4B 0.7617 0.7832 0.4864 0.3370 0.7225 0.7180 0.2791 0.1943
Table 5: Performance comparison of various LLMs. “LLaMA*” denotes Llama3-Smaug (Pal et al. 2024).
@@
The results in Figure 5 show that larger batch sizes consistently lead to better performance. This indicates that training on larger batches enables the LLM to learn stronger ranking abilities that generalize better to unseen data, which aligns with the theoretical analysis in (Lan et al. 2009). However, we also observe that larger batches make training more difficult to converge. This is intuitive, as ranking becomes more challenging with longer lists, increasing the difficulty of optimization. Therefore, how to effectively scale up listwise ranking through large-batch training or parallel optimization may be a promising direction for future research.
@@
We further investigate the impact of base LLM size on performance. We choose Qwen2.5-Instruct (Qwen et al. 2025) as our base model family due to its strong performance even at smaller scales. As shown in Figure 6, performance consistently improves as the model size increases. This suggests that our method scales effectively with larger LLMs and that the overall performance is closely tied to the capabilities of the underlying base model.
Figure 6: Performance comparison across different sizes of Qwens used as base models. The dotted line is the performance of NAIP (Zhao et al. 2025), based on 8B Llama3-Smaug (Pal et al. 2024).

### v2

@@ Introduction
Technically, the paper evaluation task aims to analyze the paper features to assess research quality from various dimensions, such as academic impact (10.1007/s11192-022-04547-8; zhao2024literature) and overall quality (lin2023automated). Existing studies mainly rely on traditional neural models or Large Language Models (LLMs) for this task:
  • Traditional methods utilize neural models, such as Multi-Layer Perceptrons (MLPs), or Long-Short Term Memory networks (LSTM), to evaluate research papers based on predefined features, including structural indicators like paper length and reference count (vergoulis2020simplifying; Ruan2020PredictingTC), as well as textual patterns (ma2021deep; yang2018automatic). However, these methods often overlook the semantic content of papers (e.g., abstract and main text), which is essential for accurate evaluation, ultimately leading to unsatisfactory performance.
  • LLM-based methods leverage rich textual information (e.g., title, abstract, and main text) to learn informative paper representations and employ a scoring module to produce evaluation scores. Empowered by the advanced semantic understanding capabilities of LLMs, these methods demonstrate strong potential in capturing the technical soundness of research papers, yielding more accurate evaluation results (lu2024ai; liu2025lmcbert; zhao2025words; de2024can).
@@
  • Reasoning-enhanced paper evaluation. Evaluating research papers requires a deep understanding of complex motivations and nuanced methodologies. This challenge is further intensified by the need to compare concurrently retrieved work. This motivates us to stimulate the reasoning mechanism of LLMs to support deep comprehension, precise comparison, and fair evaluation. While chain-of-thought reasoning (wei2022chain) offers interpretable intermediate steps, it typically requires annotated reasoning paths for supervision (weng2023large), which are scarce in paper evaluation scenarios. In contrast, latent reasoning (hao2024training) enables implicit multi-step reasoning within the hidden representations of LLMs, eliminating the need for explicit annotations. More importantly, LLM-based paper evaluation focuses on learning more informative paper representations to enhance evaluation accuracy, which aligns naturally with the latent reasoning paradigm, seamlessly integrating reasoning directly at the representation level. As such, we incorporate latent reasoning into PaperEval for comprehensive representation learning.
Despite the significant potential of latent reasoning, an effective optimization strategy is essential to guide the reasoning process toward the ultimate ranking goal of paper evaluation. Specifically, the paper evaluation task focuses on comparing and identifying valuable work within a large collection, with a primary emphasis on relative ranking rather than absolute scoring. However, learning accurate rankings is inherently more challenging, as even minor prediction errors may cause substantial shifts in the ranking positions. To address this, we propose a progressive ranking optimization strategy, which encourages the latent reasoning to progressively improve relative ranking. In particular, at each reasoning step, we compute the temperature-controlled softmax over the predicted scores of a batch of papers, which is then aligned with the ground-truth order using a listwise ranking loss. To gradually refine the LLM’s ranking, we progressively decrease the temperature during latent reasoning, making the predicted distributions increasingly sharper and more sensitive to ranking errors. This progressive refinement encourages the LLM to iteratively produce more confident and discriminative rankings within each training batch. By learning to distinguish fine-grained differences among batch samples, the LLM enhances ranking reasoning capabilities, which can naturally generalize to global ranking across the entire dataset, as theoretically supported by (10.1145/1553374.1553449).
@@
Paper evaluation aims to assess a paper’s quality or predict its academic impact. From the quality perspective, a central task is paper rating — predicting whether a paper will be accepted by peer review committees (lin2023automated). Existing methods fall into three main categories. The first leverages neural architectures like CNNs and attention-based models to capture local and global textual interactions (yang2018automatic; deng2020hierarchical). The second extracts metadata features and employs traditional models such as random forests (wang2024content). The third directly encodes the textual content using pretrained models like BERT (devlin2019bert), as in recent work (xue2023re; liu2025lmcbert). For evaluating academic impact, prior studies focus on predicting citation counts, citation levels, or other derived impact metrics (zhao2024literature), which can similarly be grouped into three strategies. Metadata-based approaches use handcrafted or extracted features with classical models like MLPs or decision trees (wang2011mining; qiu2024early; Ruan2020PredictingTC; zhang2024predicting). Graph-based methods model early citation dynamics using citation graphs and apply graph neural networks for future trend prediction (yan2024modeling; he2023h2cgl; li2023simplifying; jiang2021hints). LLM-based approaches either prompt models with a paper’s title and abstract to generate citation scores (de2024can), or map dense representations to impact-related metrics (zhao2025words). Despite these advances, accurate and fine-grained paper evaluation remains challenging. In this work, we explore how integrating retrieval techniques with the reasoning capabilities of LLMs can address this challenge.
@@
Unlike Chain-of-Thought (CoT) reasoning (wei2022chain; su2025explicit; carrow2025neural), which relies on explicitly generated intermediate steps, latent reasoning performs inference directly within a model’s hidden representations (hao2024training; biran2024hopping). This implicit approach has gained momentum in LLM-based recommendation (tang2025think; shen2025efficient; shen2025codi; liu2025lares) and retrieval tasks (ji2025learning), as it avoids the need for annotated reasoning traces while enabling richer, more informative representations. However, a key challenge remains: how to effectively supervise the latent reasoning process. On the one hand, it is essential to ensure that the reasoning process unfolds progressively toward the correct output, allowing the model to refine its prediction step by step (tang2025think; liu2025lares; ji2025learning). On the other hand, models must avoid degenerate reasoning, where the hidden states prematurely converge and hinder iterative refinement. For instance, Tang et al. (tang2025think) introduce a loss that encourages representational diversity across reasoning steps to mitigate this issue. Despite these efforts, ensuring that the model consistently refines its predictions toward accurate outcomes remains challenging. In this work, we propose a progressive optimization strategy that progressively adjusts the softmax temperature and incorporates a ranking-based loss. This design explicitly guides the latent reasoning process toward better alignment with ground-truth evaluation targets.
@@
To equip LLMs with up-to-date domain knowledge for more accurate assessment of the novelty and contributions of each target paper, we introduce a domain-aware retrieval module (depicted at the top of Figure 1). Specifically, given the title and abstract of research papers, we first employ ChatGPT (achiam2023gpt) to generate representative topic keyphrases, which are then encoded into topic embeddings using the CLIP text encoder (radford2021learning). To identify relevant work for each target paper $p_{i}$, we first compute the cosine similarity between the target topic embedding and those of all other papers in the corpus, where papers with similarity exceeding a predefined threshold $\gamma$ are retained as candidates. Considering the rapidly evolving nature of many research fields, we further filter these candidates by selecting only concurrent relevant papers, whose publication dates are closest to the target paper $p_{i}$. The resulting set, denoted as $\mathcal{R}_{i}$, contains at most $k$ papers and serves as the domain-aware reference set, providing essential contextual background to help the LLM more accurately assess the target paper’s position within the current research landscape.
@@
Inspired by ListMLE (xia2008listwise), which learns to predict rankings by maximizing the likelihood of the ground-truth order, we adapt it to the paper evaluation scenario, encouraging the predicted scores to yield a ranking consistent with the ground-truth.
@@
(1) NAID (zhao2025words): This dataset is derived from arXiv. Each paper includes its title, abstract, some metadata (e.g., paper length, number of references), and an impact score that quantifies its relative citation rank within the same domain and publication period, serving as an indicator of the paper’s future impact in its field. The dataset contains 11,118 papers in the training set and 1,237 papers in the test set.
@@
  • Traditional methods: 1) MLP-based (Ruan2020PredictingTC) method uses metadata of the target paper as input to an MLP to predict the evaluation score. We do not evaluate this method on the ICLR dataset due to the lack of metadata. 2) LSTM-based (ma2021deep) method encodes the abstract of each paper into a sequence representation and applies an LSTM network to predict the target score.
  • LLM-based methods: 3) GPT-part (de2024can) method prompts ChatGPT to predict a paper’s score based solely on its title and abstract. We use GPT-4o (hurst2024gpt) as the underlying model. 4) GPT-all (lu2024ai) approach treats the LLM as a reviewer, prompting it to read the entire paper and generate a full review, including an overall quality score. For cost considerations, we choose GPT-4o-mini as our base model. 5) SciBERT (beltagy2019scibert) method is a BERT-based model pretrained on scientific text. We fine-tune it with a simple regression module to perform paper evaluation. 6) NAIP (zhao2025words) method uses LLaMA3-Smaug (pal2024smaug) as the backbone LLM. An additional regression module is applied to the output embedding to generate the final score.
@@
  • Top-K ranking quality: Following (zhao2025words), we use NDCG@{10,20} to measure the ranking quality of the top-K recommended papers.
  • Overall ranking consistency: Following (Ng2015BetterSE), we employ Spearman’s rho and Kendall’s tau to assess how well the predicted rankings align with the ground-truth rankings.
@@
  • Pair-wise ranking: Inspired by RankNet (burges2005learning), we design a temperature-controlled pair-wise ranking loss to examine whether pair-wise supervision is more suitable for PaperEval than list-wise ranking.
@@
  • Score regression: As a control setting, similar to (zhao2025words), we remove the ranking-based objective and directly apply an MSE loss between the final predicted score and the ground-truth label. This setup allows us to examine whether ranking is a more effective learning signal than direct score supervision.
@@ Conclusion
Reproducibility Checklist
This paper:
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced. Yes.
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results. Yes.
  • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper. Yes.
Does this paper make theoretical contributions? No.
Does this paper rely on one or more datasets? Yes.
  • A motivation is given for why the experiments are conducted on the selected datasets. Yes.
  • All novel datasets introduced in this paper are included in a data appendix. Yes.
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. Yes.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. Yes.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. Yes.
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. NA.
Does this paper include computational experiments? Yes.
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. Yes.
  • Any code required for pre-processing data is included in the appendix. Yes.
  • All source code required for conducting and analyzing the experiments is included in a code appendix. Yes.
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. Yes.
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from. Yes.
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. Yes.
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. Partial.
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. Yes.
  • This paper states the number of algorithm runs used to compute each reported result. No.
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. No.
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). No.
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. Partial.
@@ Appendix A Supplementary Material
Code and Datasetshttps://github.com/ZhengWwwq/PaperEval
@@
Following NAIP (zhao2025words), we adopt LLaMA3-Smaug (pal2024smaug) as the backbone of PaperEval for fair comparison. We set the number of reasoning steps $m$ to 8 and the number of retrieved reference papers $k$ to 2. We set the temperature $\tau_{\text{max}}$ to $1.0$ and $\tau_{\text{min}}$ to $0.1$. We train the model for 5 epochs on both the NAID and ICLR datasets with a shared learning rate of $5.0\times 10^{-5}$, and select the checkpoint with the best validation performance for evaluation. All experiments are conducted on a single NVIDIA A40 GPU.
@@
Inspired by (yang2020enhancing), we also design a distribution-based similarity/ListNet (cao2007learning) loss to encourage the predicted score distribution to match the ground-truth score distribution. Specifically, we convert both the predicted scores and the ground-truth scores into soft probability distributions using a temperature-controlled softmax. The temperature varies with the reasoning step to enable coarse-to-fine comparison. The ground-truth distribution at step $j$ is computed as:
@@
We also try the RankCosine ranking loss (qin2008query) in PaperEval. RankCosine treats the predicted and ground-truth score lists as vectors and maximizes their cosine similarity, encouraging the predicted ranking to align closely with the target. In addition, we introduce temperature control to progressively sharpen the predicted vector during reasoning steps as follows:
@@
The Approximate NDCG (qin2010general) loss is a differentiable surrogate for the standard NDCG metric, enabling its direct optimization in learning-to-rank models. It overcomes the non-differentiability of the ranking operation by replacing the discrete integer rank of an item with a continuous “soft rank” $\hat{\pi}_{i}$. This soft rank is calculated from pairwise comparisons of item scores using a sigmoid function $\sigma$ scaled by a temperature parameter $\tau$.
@@
We adopt a pair-wise ranking loss inspired by RankNet (burges2005learning) to guide the model in learning relative paper quality comparisons. Specifically, at each evaluation step $t\in\{1,2,\cdots,m\}$ we focus on all pairs of papers $(i,j)$ where paper $i$ is preferred over paper $j$, means $s_{i}>s_{j}$. The model predicts scalar scores $s_{i}^{(t)}$ and $s_{j}^{(t)}$ for the two papers at step $t$, and the loss encourages $s_{i}^{(t)}$ to be higher than $s_{j}^{(t)}$ . The pair-wise loss at step $t$ is defined as:
@@
We conduct all loss design experiments on the NAID dataset. From the results in Table 4, we make two key observations: 1) ListMLE achieves the best performance among all loss functions, as its permutation probability modeling directly reflects the ground-truth ranking, making it particularly well-suited to our training objective. This also supports the theoretical analysis in (10.1145/1553374.1553449). 2) Most other listwise and pairwise losses fail to outperform simple MSE, indicating that effective ranking supervision remains challenging compared to direct score prediction.
@@
Llama* (pal2024smaug) 8B 0.9589 0.9521 0.4953 0.3438 0.7784 0.7386 0.3276 0.2285
Llama (grattafiori2024llama3herdmodels) 8B 0.9724 0.9038 0.4895 0.3405 0.7276 0.7306 0.3004 0.2091
Qwen (qwen2025qwen25technicalreport) 7B 0.9691 0.9314 0.4950 0.3446 0.7698 0.7554 0.3332 0.2320
Mistral (jiang2023mistral7b) 7B 0.8260 0.8355 0.4267 0.2944 0.7055 0.7041 0.3055 0.2141
Phi (abdin2024phi3technicalreporthighly) 4B 0.7617 0.7832 0.4864 0.3370 0.7225 0.7180 0.2791 0.1943
Table 5: Performance comparison of various LLMs. “LLaMA*” denotes Llama3-Smaug (pal2024smaug).
@@
The results in Figure 5 show that larger batch sizes consistently lead to better performance. This indicates that training on larger batches enables the LLM to learn stronger ranking abilities that generalize better to unseen data, which aligns with the theoretical analysis in (10.1145/1553374.1553449). However, we also observe that larger batches make training more difficult to converge. This is intuitive, as ranking becomes more challenging with longer lists, increasing the difficulty of optimization. Therefore, how to effectively scale up listwise ranking through large-batch training or parallel optimization may be a promising direction for future research.
@@
We further investigate the impact of base LLM size on performance. We choose Qwen2.5-Instruct (qwen2025qwen25technicalreport) as our base model family due to its strong performance even at smaller scales. As shown in Figure 6, performance consistently improves as the model size increases. This suggests that our method scales effectively with larger LLMs and that the overall performance is closely tied to the capabilities of the underlying base model.
Figure 6: Performance comparison across different sizes of Qwens used as base models. The dotted line is the performance of NAIP (zhao2025words), based on 8B Llama3-Smaug (pal2024smaug).
Effect of Temperature Decay
As illustrated in Equation (3), we employ a linearly decreasing temperature schedule. A lower temperature makes the distribution sharper, leading to a larger loss as shown in Equation (5). Formally, the gradient with respect to the scores is scaled by a factor of $\frac{1}{\tau}$. This scaling effect intensifies the model’s learning signal for refining score rankings as the temperature decreases, thereby enforcing stricter supervision on the relative orderings.
Furthermore, we explored additional temperature scheduling strategies and compared them with the scenario where no temperature decay is applied. Specifically, we implemented three alternative schedules as follows:
Exponential Decay.
$\tau^{(j)}=\tau_{\text{max}}\left(\frac{\tau_{\text{min}}}{\tau_{\text{max}}}\right)^{\frac{j}{m}}$ (17)
Cosine Decay.
$\tau^{(j)}=\tau_{\text{min}}+(\tau_{\text{max}}-\tau_{\text{min}})\cos\left(\frac{j}{m}\cdot\frac{\pi}{2}\right)$ (18)
W/o Decay.
For comparison, we also delete the temperature decay design. Instead, all temperatures are set to the same $\tau_{\text{max}}$.
Figure 7: Performance comparison across different temperature schedule designs.
From the results presented in Figure 7, we can make the following observations: 1) The exponential decay schedule causes the temperature to decrease with an increasingly flattened rate. This manifests as the model gaining an advantage in top retrieval performance, yet showing limited improvement in overall ranking quality. 2) The cosine decay schedule leads to an increasingly accelerated drop in temperature. While this results in a significant enhancement in overall ranking quality, its performance in top-K ranking quality is even inferior to that without any decay. 3) The linear decay schedule reduces the temperature at a constant rate, which is reflected in notable improvements in both top-K ranking quality and overall ranking consistency.
Computational Cost
Our setup matches NAIP, and both retrieval and latent reasoning add minimal overhead while yielding significant gains. With batch size 64, inference takes 12.3s/batch (30 GB VRAM). In PaperRec, all new papers are evaluated within 2 minutes daily, confirming the method’s efficiency.